OpenStudy (anonymous):

find x, given 1/3 log 81^(x) = -1/4

OpenStudy (zpupster):

mult by 3 both sides gives log81^x = -3/4 eliminate the log 81^x= e^-3/4 tnhen talke log base 81 bothsides leaves us with x=-3/4log(81)

OpenStudy (alekos):

thats not quite right , but nice try

OpenStudy (anonymous):

First off, multiply both sides by a constant to simplify the equation. In this equation, we will multiply both sides by 3. log^x(81)=-3/4 is what you are left with. Eliminate the exponential from the left hand side. Take the logarithm base log(81) of both sides: \[x=\frac{ i \pi (2n+1) }{ \log(\log81)) } - \frac{ \log (\frac{ 4 }{ 3 }) }{ \log(\log81)) } for n \in z\]

OpenStudy (anonymous):

Erm, guess the system for the equation showing is not working? ._.

OpenStudy (alekos):

not sure how you got this? doesnt look right either

OpenStudy (alekos):

z pupster is it log to the base 10 of 81 to the power of x?

OpenStudy (anonymous):

Anyhow........ Hold on, I'm going to show images that I have made in Microsoft word, second. Let me type the work and my final answer.

OpenStudy (alekos):

wait for zpupsters reply

OpenStudy (alekos):

sorry i meant gmk is it log to the base 10 of 81 to the power of x?

OpenStudy (alekos):

gmt, are you there?

OpenStudy (anonymous):

Here are the images, labeled with "Step 1, Step 2, etc".

OpenStudy (anonymous):

Gah, school computer renamed the files when I uploaded... but there ya go..

OpenStudy (alekos):

gmt's gone